# Fine-tuning foundational models to code diagnoses from veterinary health records

### v1

@@ Abstract
Veterinary medical records represent a large data resource for application to veterinary and One Health clinical research efforts. Use of the data is limited by interoperability challenges including inconsistent data formats and data siloing. Clinical coding using standardized medical terminologies enhances the quality of medical records and facilitates their interoperability with veterinary and human health records from other sites. Previous studies, such as DeepTag and VetTag, evaluated the application of Natural Language Processing (NLP) to automate veterinary diagnosis coding, employing long short-term memory (LSTM) and transformer models to infer a subset of Systemized Nomenclature of Medicine - Clinical Terms (SNOMED-CT) diagnosis codes from free-text clinical notes. This study expands on these efforts by incorporating all 7,739 distinct SNOMED-CT diagnosis codes recognized by the Colorado State University (CSU) Veterinary Teaching Hospital (VTH) and by leveraging the increasing availability of pre-trained large language models (LLMs). Ten freely-available pre-trained LLMs (GatorTron, MedicalAI ClinicalBERT, medAlpaca, VetBERT, PetBERT, BERT, BERT Large, RoBERTa, GPT-2, and GPT-2 XL) were fine-tuned on the free-text notes from 246,473 manually-coded veterinary patient visits included in the CSU VTH’s electronic health records (EHRs), which resulted in superior performance relative to previous efforts. The most accurate results were obtained when expansive labeled data were used to fine-tune relatively large clinical LLMs, but the study also showed that comparable results can be obtained using more limited resources and non-clinical LLMs. The results of this study contribute to the improvement of the quality of veterinary EHRs by investigating accessible methods for automated coding and support both animal and human health research by paving the way for more integrated and comprehensive health databases that span species and institutions.
Keywords: Clinical Coding, Diagnoses, Foundational Models, Large Language Models, Natural Language Processing, OMOP CDM, One Health, SNOMED-CT, Transformers, Veterinary Medical Records, Veterinary Medicine
@@ 1 Introduction
The use of veterinary medical records for clinical research efforts is often limited by interoperability challenges such as inconsistent data formats and clinical definitions, and data quality issues [dong2022automated, campbell2020computer, paynter2021veterinary, ouyang2019scoping]. Clinical coding is used to transform medical records, often in free text written by clinicians, into structured codes in a classification system like the Systemized Nomenclature of Medicine - Clinical Terms (SNOMED-CT) [chang2021use, snomedct]. SNOMED-CT is a “comprehensive clinical terminology that provides clinical content and expressivity for clinical documentation and reporting”. It is designated as a United States standard for electronic health information exchange and includes clinical findings, procedures, and observable entities for both human and non-human medicine [snomedct]. SNOMED-CT constitutes a hierarchy of standardized codes beginning with a top-level code such as “Clinical Finding” (SNOMED: 404684003) and terminating with a leaf code such as “Poisoning Due to Rattlesnake Venom” (SNOMED: 217659000). This work aims to improve methods for automating the clinical coding of veterinary medical records using hand-coded records from the Veterinary Teaching Hospital (VTH) at Colorado State University (CSU) as a training set.
@@
The coding of electronic health records (EHRs) into standardized medical terminologies allows for storage in interoperable data models such as the Observational Health Data Sciences and Informatics (OHDSI) Observational Medical Outcomes Partnership (OMOP) Common Data Model (CDM) [ohdsi]. Worldwide, more than a billion medical records from over 800 million unique patients are stored in the OMOP CDM format [ohdsiwhoweare], which is designed to promote healthcare informatics and network research. The OMOP CDM provides infrastructure that can facilitate necessary data linkages between EHRs both in veterinary and human medicine. Data linkage in medical research is a powerful tool that allows researchers to combine data from different sources related to the same patient, creating an enhanced data resource. However, several challenges hinder the full potential of its use, such as identifying accurate common entity identifiers, data inconsistencies or incompleteness, privacy concerns, and data sharing complexities [lustgarten2020veterinary].
@@
Despite these challenges, complex linkages like those involved in One Health research are a goal of veterinary and human medical institutions. One Health recognizes the interconnectedness of human, animal, and environmental health domains [cdconehealth] and the importance of considering all three to optimize health outcomes. Integrating data from different sources – such as human medical records, veterinary medical records, and environmental data – is essential for One Health research [lustgarten2020veterinary]. Application of data linkages across human, animal, and environmental data can require multiple types of data linkages; household-level linkages between humans and animals involve (1) linking specific humans (patient identification) with their companion animals (animal identification) and (2) linking the human and animal data for comparisons and analyses (human and veterinary EHRs) in a Health Insurance Portability and Accountability Act (HIPAA)-compliant manner across hospitals. For disease linkage, diagnoses need to be clearly identified across EHRs through standardized coding. This work focuses on identifying diagnoses (7,739 SNOMED-CT diagnosis codes) for animals using SNOMED-CT to pave the way for disease linkage.
@@ 1.1 Related Work
Most veterinary medical record systems do not currently include clinical coding infrastructure. The free text nature of much of the EHR, especially in veterinary medicine, makes manual clinical coding resource-intensive and difficult to scale. Reliable, supplemental, automated coding is feasible using current Natural Language Processing (NLP) methods including symbolic, knowledge-based approaches and neural network-based approaches [dong2022automated]. Symbolic AI makes use of symbols and rules to represent and model the standard practice of clinical coders. Neural networks use training data to “learn” how to match a patient’s information to the appropriate set of medical codes [dong2022automated]. Deep learning methods have been applied to clinical coding since around 2017 with the task formulated as a multi-classification problem, concept extraction problem, or Named Entity Recognition and Linking (NER + L) problem (see [durango2023named] for NER in EHRs review). How to best integrate knowledge and deep learning methods has not yet been determined and depends on the purpose of an automated clinical coding system [dong2022automated]. No matter the purpose, it is important that coders are involved from model development discussions to prospective deployment to ensure quality and usability [campbell2020computer]. Expert clinical coders can provide insight into the large coding systems and complex EHRs to help with automation. SNOMED-CT contains more than 357,000 health care concepts [snomedct], and EHRs contain many different types of documents and subsections, all of varying lengths and formats. How to best represent this varied data in an automated coding context is challenging. Research is ongoing to learn useful dense mathematical representations of patient data that accurately capture meaningful information from EHRs (e.g., [si2021deep]).
@@
Much of this ongoing work is focused on creating foundational large language models (LLMs) for different domains through learning mathematical representations of large amounts of data (see [thirunavukarasu2023large, chen2023large] for LLMs in biomedicine). These LLMs can then be fine-tuned for specific downstream tasks including clinical coding. For human medical data, Medical Information Mart for Intensive Care III (MIMIC-III) [johnson2016mimic] discharge summaries constitute a prominent publicly available foundational dataset used for benchmarking (MIMIC-IV [johnson2023mimic] has recently become available). Several prominent clinical LLMs have been trained using this data (e.g., [yang2022gatortron, huang2019clinicalbert, xie2024llama]). Other research groups have trained clinical LLMs on private, proprietary data either in addition to, or in place of, existing foundational data (e.g., [yang2022gatortron, kavuluru2015empirical, wang2023optimized, feng2024automated, van2023diagnosis, venkataraman2020fastag, chen2023meditron70b, han2023medalpaca]). No freely-available foundational datasets like MIMIC-III for the veterinary space were found. Therefore, all current veterinary clinical LLMs have been trained on private, proprietary data [zhang2019vettag, hur2020domain, farrell2023petbert]. This may be one contributing factor to the relative dearth of available veterinary clinical LLMs.
@@ 1.1.1 Motivating Work
Veterinary-specific automated clinical coding work has been done previously by other groups using a large set of manually coded veterinary medical record data from the CSU VTH [nie2018deeptag, zhang2019vettag, jiang2023vetllm]. Both DeepTag [nie2018deeptag] and VetTag [zhang2019vettag] used the text from 112,557 veterinary records hand-coded with SNOMED-CT diagnosis codes by expert coders at the CSU VTH. Over one million unlabeled clinical notes from a large private specialty veterinary group (PSVG) were additionally used for pre-training. Both models also made use of the SNOMED-CT hierarchy “to alleviate data imbalances and demonstrate such training schemes substantially benefit rare diagnoses” [zhang2019vettag]. The supplied set of target codes for each record were supplemented with each code’s hierarchical ancestry to form a hierarchical training objective. This hierarchical prediction approach “first predicts the top level codes and then sequentially predicts on a child diagnosis when its parent diagnosis is predicted to be present” [zhang2019vettag]. It was found that utilizing the hierarchy improved performance on both a held-out test set from the CSU VTH labeled data set and a test set from a private practice clinical (PP) in northern California.
@@
DeepTag (2018) is a bidirectional long-short-term memory (BiLSTM) model, which achieved impressive performance (weighted F1 score of 82.4 on CSU and 63.4 on PP data) across a set of 42 chosen SNOMED-CT diagnosis codes. VetTag (2019) expanded on the original concept of DeepTag by employing a transformer [vaswani2017transformer] model architecture and expanding the size of the set of chosen SNOMED-CT codes to 4,577. Pre-training was performed using the PSVG data and fine-tuning was done on the CSU VTH data. The expanded scope and improved design of VetTag made it the state-of-the-art in the automated veterinary diagnosis coding field (weighted F1 score of 66.2 on CSU VTH and 48.6 on PP data). VetLLM (2023) [jiang2023vetllm] evaluated a recently developed foundational LLM, Alpaca-7B [alpaca] in a zero-shot setting, supplying it 5,000 veterinary notes to identify nine high-level SNOMED-CT disease codes (weighted F1 score of 74.7 on CSU VTH and 63.7 on PP data). The DeepTag, VetTag, and VetLLM trained models and data are not publicly available due to use of proprietary medical data.
1.1.2 Human Clinical LLMs
@@
FasTag [venkataraman2020fastag], a study into the cross-applicability of human and veterinary clinical data, showed that applying human medical data to veterinary clinical tasks has promise considering the limited availability of veterinary LLMs. Building on FasTag, three prominent, publicly available human medicine LLMs that can be fine-tuned on a variety of downstream tasks including clinical coding were considered and tested for this study: GatorTron (2022) [yang2022gatortron], MedicalAI ClinicalBERT (2023) [wang2023optimized], and medAlpaca (2023) [han2023medalpaca]. Others that were not specifically considered for this study due to computational limitations include MEDITRON (2023) [chen2023meditron70b] and Me-LLaMA (2024) [xie2024llama].
@@
GatorTron is one of the largest clinical language models in the world and was created through a partnership between the University of Florida (UF) and NVIDIA. The model was pre-trained on a corpus of >90 billion words of clinically-adjacent text from de-identified UF human clinical notes (2.9 million clinical notes from 2.4 million patients), the publicly-available MIMIC-III dataset, PubMed, and Wikipedia. The model displayed strong improvements, relative to previous models, in five downstream NLP tasks including “clinical concept extraction, medical relation extraction, semantic textual similarity, natural language inference, and medical question answering” [yang2022gatortron]. GatorTron is a versatile LLM that can be trained for many different tasks, including veterinary clinical coding.
@@
MedicalAI ClinicalBERT is a clinical language model constructed by researchers at Fudan University. It was selected for exploration as a lighter-weight alternative to the massive GatorTron. It should be noted that other models with the name “ClinicalBERT” [huang2019clinicalbert] exist, but to encompass as much data and prior work as possible, a relevant fine-tuned ClinicalBERT model, MedicalAI, was used for this study. The model was pre-trained on 1.2 billion words from clinical records representing a diverse set of diseases. It was then fine-tuned on clinical data from hospitalized patients with Type 2 Diabetes (T2D) who received insulin therapy from January 2013 to April 2021 in the Department of Endocrinology and Metabolism, Zhongshan Hospital and Qingpu Hospital, in Shanghai, China. It was benchmarked on an extraction of 40 symptom labels from clinical notes including free-text descriptions of present illness and physical examination [wang2023optimized], similar to the task of this study.
@@
medAlpaca was inspired by the potential for LLMs like those in the GPT series to “[improve] medical workflows, diagnostics, patient care, and education” [han2023medalpaca]. It was meant to be a resource that could be locally deployed to protect patient privacy. A dataset named Medical Meadow, aimed at encapsulating a broad set of clinical use-cases, was curated and used to fine-tune the LLaMA [touvron2023llama] foundational model. Medical Meadow consisted of question-answer pairs from Anki Medical Curriculum flashcards, medically-relevant Stack Exchange forums, Wikidoc articles, and various foundational clinical datasets. medAlpaca was benchmarked in a zero-shot setting against portions of the United States Medical Licensing Examination (USMLE) [han2023medalpaca].
1.1.3 Veterinary Clinical LLMs

### v2

@@ Abstract
Veterinary medical records represent a large data resource for application to veterinary and One Health clinical research efforts. Use of the data is limited by interoperability challenges including inconsistent data formats and data siloing. Clinical coding using standardized medical terminologies enhances the quality of medical records and facilitates their interoperability with veterinary and human health records from other sites. Previous studies, such as DeepTag and VetTag, evaluated the application of Natural Language Processing (NLP) to automate veterinary diagnosis coding, employing long short-term memory (LSTM) and transformer models to infer a subset of Systemized Nomenclature of Medicine - Clinical Terms (SNOMED-CT) diagnosis codes from free-text clinical notes. This study expands on these efforts by incorporating all 7,739 distinct SNOMED-CT diagnosis codes recognized by the Colorado State University (CSU) Veterinary Teaching Hospital (VTH) and by leveraging the increasing availability of pre-trained language models (LMs). 13 freely-available pre-trained LMs (GatorTron, MedicalAI ClinicalBERT, medAlpaca, VetBERT, PetBERT, BERT, BERT Large, RoBERTa, GPT-2, GPT-2 XL, DeBERTa V3, ModernBERT, and Clinical ModernBERT) were fine-tuned on the free-text notes from 246,473 manually-coded veterinary patient visits included in the CSU VTH’s electronic health records (EHRs), which resulted in superior performance relative to previous efforts. The most accurate results were obtained when expansive labeled data were used to fine-tune relatively large clinical LMs, but the study also showed that comparable results can be obtained using more limited resources and non-clinical LMs. The results of this study contribute to the improvement of the quality of veterinary EHRs by investigating accessible methods for automated coding and support both animal and human health research by paving the way for more integrated and comprehensive health databases that span species and institutions.
Keywords: Clinical Coding, Diagnoses, Foundational Models, Language Models, Natural Language Processing, OMOP CDM, One Health, SNOMED-CT, Transformers, Veterinary Medical Records, Veterinary Medicine
Author Summary
In this study, we explored the use of advanced natural language processing (NLP) techniques to improve the quality and interoperability of veterinary medical records. By leveraging a variety of pre-trained language models (LMs) and a labeled training dataset curated by expert medical coders to apply standardized medical terminologies to diagnoses from free-text clinical notes, we demonstrate a powerful use-case for recent developments in NLP technologies. Our findings suggest that complex LMs fine-tuned on large volumes of curated data yield best results for quick and reliable automated diagnosis coding. However, we also show that comparable results can be attained using a more minimal set of computational and data resources. We believe this study can provide guidance for other clinical sites interested in enhancing the quality of electronic health records in both the veterinary and human domains. Accurate, automated medical record coding methods may facilitate and encourage clinical research and data sharing in the veterinary, human, and One Health contexts.
@@ 1 Introduction
The use of veterinary medical records for clinical research efforts is often limited by interoperability challenges such as inconsistent data formats and clinical definitions, and data quality issues [69, 63, 100, 99]. Clinical coding is used to transform medical records, often in free text written by clinicians, into structured codes in a classification system like the Systemized Nomenclature of Medicine - Clinical Terms (SNOMED-CT) [64, 96]. SNOMED-CT is a “comprehensive clinical terminology that provides clinical content and expressivity for clinical documentation and reporting”. It is designated as a United States standard for electronic health information exchange and includes clinical findings, procedures, and observable entities for both human and non-human medicine [96]. SNOMED-CT constitutes a hierarchy of standardized codes beginning with a top-level code such as “Clinical Finding” (SNOMED: 404684003) and terminating with a leaf code such as “Poisoning Due to Rattlesnake Venom” (SNOMED: 217659000). This work aims to improve methods for automating the clinical coding of veterinary medical records using hand-coded records from the Veterinary Teaching Hospital (VTH) at Colorado State University (CSU) as a training set.
@@
The coding of electronic health records (EHRs) into standardized medical terminologies allows for storage in interoperable data models such as the Observational Health Data Sciences and Informatics (OHDSI) Observational Medical Outcomes Partnership (OMOP) Common Data Model (CDM) [95]. Worldwide, more than a billion medical records from over 800 million unique patients are stored in the OMOP CDM format [118], which is designed to promote healthcare informatics and network research. The OMOP CDM provides infrastructure that can facilitate necessary data linkages between EHRs both in veterinary and human medicine. Data linkage in medical research is a powerful tool that allows researchers to combine data from different sources related to the same patient, creating an enhanced data resource. However, several challenges hinder the full potential of its use, such as identifying accurate common entity identifiers, data inconsistencies or incompleteness, privacy concerns, and data sharing complexities [89].
@@
Despite these challenges, complex linkages like those involved in One Health research are a goal of veterinary and human medical institutions. One Health recognizes the interconnectedness of human, animal, and environmental health domains [73] and the importance of considering all three to optimize health outcomes. Integrating data from different sources – such as human medical records, veterinary medical records, and environmental data – is essential for One Health research [89]. Application of data linkages across human, animal, and environmental data can require multiple types of data linkages; household-level linkages between humans and animals involve (1) linking specific humans (patient identification) with their companion animals (animal identification) and (2) linking the human and animal data for comparisons and analyses (human and veterinary EHRs) in a Health Insurance Portability and Accountability Act (HIPAA)-compliant manner across hospitals. For disease linkage, diagnoses need to be clearly identified across EHRs through standardized coding. This work focuses on identifying diagnoses (7,739 SNOMED-CT diagnosis codes) for animals using SNOMED-CT to pave the way for disease linkage.
@@ 1.1 Related Work
Most veterinary medical record systems do not currently include clinical coding infrastructure. The free text nature of much of the EHR, especially in veterinary medicine, makes manual clinical coding resource-intensive and difficult to scale. Reliable, supplemental, automated coding is feasible using current Natural Language Processing (NLP) methods including symbolic, knowledge-based approaches and neural network-based approaches [69, 90, 67]. Symbolic AI makes use of symbols and rules to represent and model the standard practice of clinical coders. Neural networks use training data to “learn” how to match a patient’s information to the appropriate set of medical codes [69]. Deep learning methods have been applied to clinical coding since around 2017 with the task formulated as a multi-classification problem, concept extraction problem, or Named Entity Recognition and Linking (NER + L) problem (see [70] for NER in EHRs review). How to best integrate knowledge and deep learning methods has not yet been determined and depends on the purpose of an automated clinical coding system [69]. No matter the purpose, it is important that coders are involved from model development discussions to prospective deployment to ensure quality and usability [63]. Expert clinical coders can provide insight into the large coding systems and complex EHRs to help with automation. SNOMED-CT contains more than 357,000 health care concepts [96], and EHRs contain many different types of documents and subsections, all of varying lengths and formats. How to best represent this varied data in an automated coding context is challenging. Research is ongoing to learn useful dense mathematical representations of patient data that accurately capture meaningful information from EHRs (e.g., [105]).
@@
Much of this ongoing work is focused on creating foundational language models (LMs) for different domains through learning mathematical representations of large amounts of data (see [109, 65] for LMs in biomedicine). These LMs can then be fine-tuned for specific downstream tasks including clinical coding. For human medical data, Medical Information Mart for Intensive Care III (MIMIC-III) [83] discharge summaries constitute a prominent publicly available foundational dataset used for benchmarking (MIMIC-IV [82] has recently become available). Several prominent clinical LMs have been trained using this data (e.g., [120, 78, 119]). Other research groups have trained clinical LMs on private, proprietary data either in addition to, or in place of, existing foundational data (e.g., [120, 84, 116, 72, 112, 114, 66, 75]). No freely-available foundational datasets like MIMIC-III for the veterinary space were found. Therefore, all current veterinary clinical LMs have been trained on private, proprietary data [121, 79, 71]. This may be one contributing factor to the relative dearth of available veterinary clinical LMs.
@@ 1.1.1 Motivating Work
Veterinary-specific automated clinical coding work has been done previously by other groups using a large set of manually coded veterinary medical record data from the CSU VTH [94, 121, 81]. Both DeepTag [94] and VetTag [121] used the text from 112,557 veterinary records hand-coded with SNOMED-CT diagnosis codes by expert coders at the CSU VTH. Over one million unlabeled clinical notes from a large private specialty veterinary group (PSVG) were additionally used for pre-training. Both models also made use of the SNOMED-CT hierarchy “to alleviate data imbalances and demonstrate such training schemes substantially benefit rare diagnoses” [121]. The supplied set of target codes for each record were supplemented with each code’s hierarchical ancestry to form a hierarchical training objective. This hierarchical prediction approach “first predicts the top level codes and then sequentially predicts on a child diagnosis when its parent diagnosis is predicted to be present” [121]. It was found that utilizing the hierarchy improved performance on both a held-out test set from the CSU VTH labeled data set and a test set from a private practice clinic (PP) in northern California.
@@
DeepTag (2018) is a bidirectional long-short-term memory (BiLSTM) model, which achieved impressive performance (weighted F1 score of 82.4 on CSU and 63.4 on PP data) across a set of 42 chosen SNOMED-CT diagnosis codes. VetTag (2019) expanded on the original concept of DeepTag by employing a transformer [113] model architecture and expanding the size of the set of chosen SNOMED-CT codes to 4,577. Pre-training was performed using the PSVG data and fine-tuning was done on the CSU VTH data. The expanded scope and improved design of VetTag made it the state-of-the-art in the automated veterinary diagnosis coding field (weighted F1 score of 66.2 on CSU VTH and 48.6 on PP data). VetLLM (2023) [81] evaluated a recently developed foundational LM, Alpaca-7B [108] in a zero-shot setting, supplying it 5,000 veterinary notes to identify nine high-level SNOMED-CT disease codes (weighted F1 score of 74.7 on CSU VTH and 63.7 on PP data). The DeepTag, VetTag, and VetLLM trained models and data are not publicly available due to use of proprietary medical data.
1.1.2 Human Clinical LMs
@@
FasTag [114], a study into the cross-applicability of human and veterinary clinical data, showed that applying human medical data to veterinary clinical tasks has promise considering the limited availability of veterinary LMs. Building on FasTag, four prominent, publicly available human medicine LMs that can be fine-tuned on a variety of downstream tasks including clinical coding were considered and tested for this study: GatorTron (2022) [120], MedicalAI ClinicalBERT (2023) [116], medAlpaca (2023) [75], and Clinical ModernBERT (2025) [85]. Others that were not specifically considered for this study due to computational limitations include MEDITRON (2023) [66] and Me-LLaMA (2024) [119].
@@
GatorTron is one of the largest clinical language models in the world and was created through a partnership between the University of Florida (UF) and NVIDIA. The model was pre-trained on a corpus of >90 billion words of clinically-adjacent text from de-identified UF human clinical notes (2.9 million clinical notes from 2.4 million patients), the publicly-available MIMIC-III dataset, PubMed, and Wikipedia. The model displayed strong improvements, relative to previous models, in five downstream NLP tasks including “clinical concept extraction, medical relation extraction, semantic textual similarity, natural language inference, and medical question answering” [120]. GatorTron is a versatile LM that can be trained for many different tasks, including veterinary clinical coding.
@@
MedicalAI ClinicalBERT is a clinical language model constructed by researchers at Fudan University. It was selected for exploration as a lighter-weight alternative to the massive GatorTron. It should be noted that other models with the name “ClinicalBERT” [78] exist, but to encompass as much data and prior work as possible, a relevant fine-tuned ClinicalBERT model, MedicalAI, was used for this study. The model was pre-trained on 1.2 billion words from clinical records representing a diverse set of diseases. It was then fine-tuned on clinical data from hospitalized patients with Type 2 Diabetes (T2D) who received insulin therapy from January 2013 to April 2021 in the Department of Endocrinology and Metabolism, Zhongshan Hospital and Qingpu Hospital, in Shanghai, China. It was benchmarked on an extraction of 40 symptom labels from clinical notes including free-text descriptions of present illness and physical examination [116], similar to the task of this study.
@@
medAlpaca was inspired by the potential for LMs like those in the GPT series to “[improve] medical workflows, diagnostics, patient care, and education” [75]. It was meant to be a resource that could be locally deployed to protect patient privacy. A dataset named Medical Meadow, aimed at encapsulating a broad set of clinical use-cases, was curated and used to fine-tune the LLaMA [110] foundational model. Medical Meadow consisted of question-answer pairs from Anki Medical Curriculum flashcards, medically-relevant Stack Exchange forums, Wikidoc articles, and various foundational clinical datasets. medAlpaca was benchmarked in a zero-shot setting against portions of the United States Medical Licensing Examination (USMLE) [75].
Clinical ModernBERT [85] is a clinical language model based on the ModernBERT [117] architecture. ModernBERT utilizes rotary positional embeddings (RoPE), unpadding, and flash attention to improve model efficiency. The model was pretrained on over 13 billion tokens from the publicly-available MIMIC-IV dataset, PubMed, and ICD-9 through ICD-12 concept descriptions. Clinical ModernBERT outperformed other BERT-based models on EHR, PubMed-NCT, and MedNER classification benchmark tasks.
1.1.3 Veterinary Clinical LMs